A PARTICLE MOVES ALONG A STRAIGHT LINE SUCH THAT AT TIME T ITS DISPLACEMENT FROM AFIXED PINT O ON THE LINE IS 3T SQUARE – 2 WHOLE POWER M. THE VELOCITY OF THE PARTICLE WHEN T=2S IS

Grade:9

1 Answers

Karan Gupta
4 years ago
s =(3t^2 - 2)^m v = ds/dt = m(3t^2 -2)^m^-^1(16t) v = m(3(2)^2 -2)^m^-^1(12) =12m(10)^m^-^1ms^-^2since velocity is the first derivitive of displacecemt with respect to time we differentiate the function of displacecemt to get the expression for velocity and plug in the given time in the expression to get the answer.